OpenStudy (anonymous):

Arianna kicks a soccer ball off the ground and into the air with an initial velocity of 42 feet per second. Assume the starting height of the ball is 0 feet. Approximately what maximum height does the soccer ball reach?

OpenStudy (anonymous):

You can use the kinematics equation in this form. v^2 = u^2 + 2(a)(s) where v represents final velocity and u represents initial velocity. a is acceleration and s is distance. Substitute in all the numbers, 0 = 42^2 + 2 ( -9.81) (s) 19.62s =1764 s=89.9 feet